Solve each equation or system of equations.
y=x-3
7x-2y=21

Answer by checkley75(3666)

y=x-3 then
7x-2(x-3)=21
7x-2x+6=21
5x=21-6
5x=15
x=15/5
x=3 answer
y=5+3
y=8 answer
proof
7*3-2(3-3)=21
21-2*0=21
21=21